OpenStudy (anonymous):

MEDALSS Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9 an = -579 + 42(n + 1) an = -579 + 42(n - 1) an = -579 - 42(n + 1) an = -579 - 42(n - 1)

OpenStudy (anonymous):

it is b

OpenStudy (anonymous):

because \[a _{n}=a _{1}+(n-1)d\]

OpenStudy (anonymous):

@miss.sweetiepie Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12, ... an = -15 + 9(n + 1) an = -15 x 9(n - 1) an = -15 + 9(n + 2) an = -15 + 9(n - 1)

OpenStudy (anonymous):

D be because of the equation I mentioned before

OpenStudy (anonymous):

could you testimonial and fan?

OpenStudy (anonymous):

@Miss.SweetiePie fan!!

OpenStudy (anonymous):

Thank you! Do you need help in anything else